OpenStudy (anonymous):

Write a problem involving the addition or multiplication of two integers with different signs. Be sure to include your answer to this problem

OpenStudy (texaschic101):

3 x -6 = - 18 -4 + 8 = 4

OpenStudy (anonymous):

22[4 + (3 - 6)2] - 13